OpenStudy (anonymous):

two integers that have a product of 8 and sum of -2 what are the integers

OpenStudy (kc_kennylau):

xy=8; x+y=-2

OpenStudy (kc_kennylau):

Does not exist; perhaps you mean -8 as the product?

OpenStudy (anonymous):

i think it was 2 as the sum srry

OpenStudy (anonymous):

You still get a non real solution

OpenStudy (kc_kennylau):

xy=8; x+y=2

OpenStudy (kc_kennylau):

Still does not exist

OpenStudy (kc_kennylau):

Did you mean product=-8 and sum=2

OpenStudy (anonymous):

yes srry

OpenStudy (kc_kennylau):

4*-2=-8 4+(-2)=2

OpenStudy (anonymous):

Answer is 4 and -2

OpenStudy (anonymous):

thank you both

OpenStudy (kc_kennylau):

no problem